OpenStudy (abbydale74):

Which expression is equivalent to seven to the third divided by seven to the fifth?

OpenStudy (alekos):

7^3/7^5 = 7^(3-5) = 7^2

OpenStudy (alekos):

sorry thats 7^(-2)

OpenStudy (abbydale74):

thank you!